a. Use the factor three to determine the prime factors and prime products of 1260



b. Use vertical and horizontal algorithms to find the difference of 709-568. Explain the borrow concepts.

1
Expert's answer
2022-03-07T13:28:01-0500

Part(a)

Factors of 1260

Factors of 1260 are integers that can be divided evenly into 1260. It has total 36 factors of which 1260 is the biggest factor and the prime factors of 1260 are 2, 3, 5, 7. The sum of all factors of 1260 is 4368.

  • All Factors of 1260: 1, 2, 3, 4, 5, 6, 7, 9, 10, 12, 14, 15, 18, 20, 21, 28, 30, 35, 36, 42, 45, 60, 63, 70, 84, 90, 105, 126, 140, 180, 210, 252, 315, 420, 630 and 1260
  • Prime Factors of 1260: 2, 3, 5, 7 
  • Prime Factorization of 1260: 22 × 32 × 51 × 71
  • Sum of Factors of 1260: 4368



Part(b)


709-568

Horizontal algorithm-

709=700+0+9

568=500+60+8


749-568=(700-500)+(0-60)+(9-8)=200-60+1=141


Vertical algorithm-


"7|0^1|9"

"-5|6|8"

"\overline{ 1|4|1}"


Here we borrow a carry from 7 so that we can perform subtraction.